OpenStudy (anonymous):

5 _________ 4 + sqrt (2

OpenStudy (anonymous):

Is your question to rationalise this? \[\frac{ 5 }{ 4+ \sqrt{2} }\]

OpenStudy (anonymous):

to simplify

OpenStudy (anonymous):

Rationalising the denominator will get the fraciton in a simpler form. Do you know how to do this or would you like me to explain?

OpenStudy (anonymous):

i take the denominator and multiply it to the top and the bottom . right ??

OpenStudy (anonymous):

Almost, but you have to remember to flip the sign of the irrational part. So for example, if your fraction was: \[\frac{ 2 }{3 + \sqrt{5} }\] Then you would multiply top and bottom by: \[3-\sqrt{5}\] Does this look familiar? Are you able to apply this to your question?

OpenStudy (anonymous):

is this it